# On the volume of K-semistable Fano manifolds

### v1

@@ 4. Volume estimates and minimal rational curves
4.1. Proof of Theorem 1.4 and Theorem 1.1
The calculation in the previous section is based on the test of K-stability via standard blow-ups. In this section, we prove Theorem 1.4 and Theorem 1.1 by considering the test of K-stability via weighted blow-ups along (Zariski open sets of) minimal rational curves.
Proof of Theorem 1.4.
Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ be minimal rational curve, or equivalently an immersed rational curve that satisfies $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ with $12\leq d\leq n+$ 1 and $d=f^{*}(-K_{X})\cdot{\mathbb{P}}^{1}=(-K_{X})\cdot f({\mathbb{P}}^{1})$. The (relative) normal bundle has the splitting:
$N_{f/X}=f^{*}TX/T{\mathbb{P}}^{1}={\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
@@
Set $Z=f({\mathbb{P}}^{1})$ which is an irreducible curve with possible nodal singularities. Fix a regular point $x\in Z$, there exists a Zariski open neighborhood $U$ and coordinates $\{z_{i};i=1,\dots,n\}$ such that
@@
    For any $21\leq i\leq d-$ 2, ${\rm d}z_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}$ over $Z\cap U$;
@@
    For any $d-1\leq j\leq n-1$, ${\rm d}z_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}$ over $Z\cap U$.
For any $\ell\in{\mathbb{Z}}_{>0}$, we can define a divisorial valuation $v={\rm ord}_{E}$: for any $f\in{\mathcal{O}}(U)$,
(17) $v(f)={\rm min}\{|I|+\ell|J|;f=\sum_{I,J}{a_{IJ}}z^{\prime I}z^{\prime\prime J},a_{IJ}=a_{IJ}(z_{n})\neq 0\}$
@@
where $z^{\prime}=\{z_{1},\dots,z_{d-2}\}$ (resp. $z^{\prime\prime}=\{z_{d-1},\cdots,z_{n-1}\}$), and $z^{\prime I}=z_{1}^{i_{1}}\cdots z_{d-2}^{i_{d-2}}$ for $I=(i_{1},\dots,i_{d-2})\in{\mathbb{N}}^{d-2}$ with $|I|=i_{1}+\cdots+i_{d-2}$ (resp. $z^{\prime\prime J}=z_{d-1}^{j_{1}}\cdots z_{n-1}^{j_{n-d+1}}$ for $J=(j_{1},\cdots,j_{n-d+1})\in{\mathbb{N}}^{n-d+1}$ with $|J|=j_{1}+\cdots+j_{n-d+1}$).
@@
When restricted to $U$, the valuation $v$ corresponds to the exceptional divisor $E$ of the weighted blowup of $X\cap U$ along $Z\cap U$ with weights $(\underbrace{1,\dots,1}_{d-2},\underbrace{\ell,\cdots,\ell}_{n-d+1})$ (See [QR12] for more a general set-up of weighted blowups). Set ${\mathcal{I}}_{x}={\mathcal{I}}_{x}({\rm ord}_{E})=\{f\in{\mathcal{O}}_{X};v(f)\geq x\}$. Then we have the exact sequence:
@@
For any $[s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$, we know that with respect to a local trivialization of $L^{k}$, the section $s$ is represented locally over $U$ by a holomorphic function with the following expression modulo terms with $|I|+\ell|J|>y$,
$s$ $=$ $\sum_{|I|+\ell|J|=y}a_{IJ}z^{\prime I}z^{\prime\prime J}=\sum_{k=0}^{\lfloor y/\ell\rfloor}\sum_{|I|=y-\ell k,|J|=k}a_{IJ}z^{\prime I}z^{\prime\prime J}$

### v2

@@ 4. Volume estimates and minimal rational curves
The calculation in the previous section is based on the test of K-stability via standard blow-ups. In this section, we prove Theorem 1.1 by carrying out a test of K-stability via a particular weighted blow-ups along minimal rational curves (see Example (4.16) for the motivating example). We also need to overcome the difficulty of possible singularities on the rational curve.
4.1. Two divisorial valuations: $v_{1}$ and $v_{2}$
Let $f:{\mathbb{P}}^{1}\rightarrow X$ be a minimal rational curve. In other words, it is an immersed rational curve that satisfies $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ with $2\leq d\leq n+1$ and $d=f^{*}(-K_{X})\cdot{\mathbb{P}}^{1}=(-K_{X})\cdot f({\mathbb{P}}^{1})$. The (relative) normal bundle has the splitting:
(12) $N_{f/X}=f^{*}TX/T{\mathbb{P}}^{1}={\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}.$
@@ 4.1. Two divisorial valuations: $v_{1}$ and $v_{2}$
Set $Z=f({\mathbb{P}}^{1})$ which is an irreducible curve with possibly a finite number of singularities $Z^{\mathrm{sing}}=\{p_{1},\dots,p_{k}\}$ and $f:{\mathbb{P}}^{1}\rightarrow Z$ is the normalization of $Z$. Set $f^{-1}(p_{i})=\{t_{i}^{j};1\leq j\leq m_{i}\}$ to be the inverse image of the singularity $p_{i}$ under $f$ so that $f^{-1}(Z^{\mathrm{sing}})=\cup_{i}f^{-1}(p_{i})$.
For any point $t\in{\mathbb{P}}^{1}$, we first construct a complex analytic coordinate chart near $f(t)\in X$ in the following way, which is adapted to the splitting (12). Choose a disk $B_{t}:=B_{t}(\epsilon)\subset{\mathbb{P}}^{1}$ of radius $\epsilon>0$ (with respect to the round spherical metric) centered at $t$, such that $f:B_{t}\rightarrow X$ is an embedding. Since $f:{\mathbb{P}}^{1}\rightarrow X$ is an immersion and $f^{-1}(Z^{\mathrm{sing}})$ is a finite set, it is easy to see that we can choose $\epsilon$ sufficiently small such that $B_{t}$ exists for all $t\in{\mathbb{P}}^{1}$. Moreover, by choosing a uniform $0<\epsilon\ll 1$, we can choose an open neighborhood $U=U_{t}$ of $f(t)$ and complex analytic coordinates ${\mathbf{z}}={\mathbf{z}}_{t}:=\{z_{1},\dots,z_{n}\}$ on $U$ such that
@@
    $f(B_{t})\cap U_{t}=\{z_{1}=\cdots=z_{n-1}=0\}$;
@@
    For any $1\leq i\leq d-2$, $f^{*}{\rm d}z_{i}$ is a local generator of the $i$-th summand of ${\mathcal{O}}(-1)^{\oplus(d-2)}\hookrightarrow N^{\vee}_{f/X}$ over $B_{t}$;
  4. (4)
    For any $d-1\leq j\leq n-1$, $f^{*}{\rm d}z_{j}$ is a local generator of the $j$-th summand of ${\mathcal{O}}^{\oplus(n-d+1)}\hookrightarrow N^{\vee}_{f/X}$ over $B_{t}$.
We now define a divisorial valuation $v={\rm ord}_{E}$ whose center over $X$ is $Z$. Pick any rational function $\psi$ on $X$. We assume that $\psi$ is defined on a Zariski open set $W$ such that $W\cap Z\neq\emptyset$. Pick any point $t\in f^{-1}(W\cap Z)$. The restriction $\psi$ to $U=U_{t}$ becomes holomorphic near $0\in U$. Expand the restriction $f|U$ near $0\in U$:
(13) $\psi({\mathbf{z}})=\sum_{I,J}{b_{IJ}}(z_{n}){\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J}$
@@
where ${\mathbf{z}}^{\prime}=\{z_{1},\dots,z_{d-2}\}$ (resp. ${\mathbf{z}}^{\prime\prime}=\{z_{d-1},\cdots,z_{n-1}\}$), and ${\mathbf{z}}^{\prime I}=z_{1}^{i_{1}}\cdots z_{d-2}^{i_{d-2}}$ for $I=(i_{1},\dots,i_{d-2})\in{\mathbb{N}}^{d-2}$ with $|I|=i_{1}+\cdots+i_{d-2}$ (resp. ${\mathbf{z}}^{\prime\prime J}=z_{d-1}^{j_{1}}\cdots z_{n-1}^{j_{n-d+1}}$ for $J=(j_{1},\cdots,j_{n-d+1})\in{\mathbb{N}}^{n-d+1}$ with $|J|=j_{1}+\cdots+j_{n-d+1}$). We then define: for $\ell=1$ or $\ell=2$
(14) $v_{\ell}(\psi)={\rm min}\{|I|+\ell|J|;\psi=\sum_{I,J}{b_{IJ}}{\mathbf{z}}^{\prime I}{\mathbf{z}}^{\prime\prime J},b_{IJ}=b_{IJ}(z_{n})\not\equiv 0\}.$
We claim that this is independent of $t\in f^{-1}(W\cap Z)$. Since $Z\setminus W$ is a finite set, $W\cap Z$ is connected. So we just need to verify the claim locally. Assume that $t^{\prime}$ is near $t$ so that $x^{\prime}=f(t^{\prime})$ is near $x=f(t)$ and we can find a sequence of points $\{t_{i};i=1,\dots,N\}$ and small discs $B_{t_{i}}(\epsilon_{1})\subset B_{t_{i}}(\epsilon)=B_{t_{i}}$ with $1<\epsilon_{1}\ll\epsilon$ such that $B_{t_{0}}(\epsilon_{1})=B_{t}(\epsilon_{1})$, $B_{t_{N}}(\epsilon_{1})=B_{t^{\prime}}(\epsilon_{1})$ and $B_{t_{i}}(\epsilon_{1})\cap B_{t_{i+1}}(\epsilon_{1})\supset B_{t_{i}}(\epsilon_{1}/3)\cup B_{t_{i+1}}(\epsilon_{1}/3)$. In particular, this implies such that $U_{t_{i}}\cap U_{t_{i+1}}\neq\emptyset$. For each $1\leq i\leq N-1$, the transition function on $U_{t_{i}}\cap U_{t_{i+1}}$ from $(U_{t_{i}},{\mathbf{z}}={\mathbf{z}}_{t_{i}})$ and $(U_{t_{i+1}},\hat{{\mathbf{z}}}:={\mathbf{z}}_{t_{i+1}})$ is of the form:
(15) $z_{k}$ $=$ $a_{k}(\hat{z}_{n})\hat{z}_{k}+R_{2},1\leq k\leq n-1,\quad z_{n}=g_{n}(\hat{z}_{n})+R_{1}$
where $a_{k}(\hat{z}_{n})$ is a non-zero holomorphic transition function of ${\mathcal{O}}_{{\mathbb{P}}^{1}}(-1)$ for $1\leq k\leq d-2$ and of ${\mathcal{O}}_{{\mathbb{P}}^{1}}$ for $d-1\leq k\leq n-1$, $g_{n}(\hat{z}_{n})$ is an invertible holomorphic function and $R_{2}$ (resp. $R_{1}$) are holomorphic functions whose expansion consists only of terms of degree at least 2 (resp. at least 1) in $\hat{z}_{1},\dots,\hat{z}_{n-1}$. Assume $\psi|_{U_{t_{i}}}$ has the expansion (13) centered at $t_{i}\in U_{t_{i}}$ with all terms of weighted degree $|I|+\ell|J|$ at least $y$, i.e. the right-hand-side of (14) is equal to $y$. Then
$\psi({\mathbf{z}})$ $=$ $\sum_{|I|+\ell|J|\geq y}b_{IJ}(g_{n}(\hat{z}_{n})+R_{1})(\mathbf{a}^{\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime}+R_{2})^{I}(\mathbf{a}^{\prime\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime\prime}+R_{2})^{J}$
$=$ $\sum_{|I|+\ell|J|=y}b_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime})^{I}(\mathbf{a}^{\prime\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime\prime})^{J}+\sum_{|I|+\ell|J|=y}b^{t}_{IJ}(g_{n}(\hat{z}_{n}))(\mathbf{a}^{\prime}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime})^{I}R^{\prime J}_{2}+R^{w}_{y+1}$
$=$ $\sum_{K,L}\hat{b}_{KL}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime K}\hat{{\mathbf{z}}}^{\prime\prime L}$
where $(\mathbf{a}^{\prime}\hat{\mathbf{z}}^{\prime})^{I}=\prod_{k=1}^{d-2}a_{k}^{i_{k}}z_{k}^{i_{k}}$, $(\mathbf{a}^{\prime\prime}\hat{\mathbf{z}}^{\prime\prime})^{J}=\prod_{k=d-1}^{n-1}a_{k}^{j_{k}}z_{k}^{j_{k}}$, $R^{\prime}_{2}$ consists of terms of degree at least $2$ in $(\hat{z}_{1},\dots,\hat{z}_{d-2})$, $R^{w}_{y+1}$ consists of terms of weighted degree $|I|+\ell|J|$ at least $y+1$ and the last expression is the expansion of $f|U_{t_{i+1}}$ near the center $t_{i+1}\in U_{t_{i+1}}$. It is then easy to see that for $\ell=1$ or $\ell=2$,
${\rm min}\{|K|+\ell|L|;\psi=\sum_{K,L}\hat{b}_{KL}(\hat{z}_{n})\hat{{\mathbf{z}}}^{\prime K}\hat{{\mathbf{z}}}^{\prime\prime L},\hat{b}_{KL}\not\equiv 0\}$
is at least the quantity on the right-hand-side of (14). Since we can switch the role of $t_{i}$ and $t_{i+1}$, we know that the two quantities are actually equal. By passing from $t_{0}=t$ to $t_{N}=t^{\prime}$, we conclude that for $\ell=1,2$, $v_{\ell}(f)$ does not depend on $t$ and $t^{\prime}$.
Remark 4.1.
Definition of quasi-monomial valuations by using analytic expansions under complex analytic coordinates has also recently appeared in [MP25, 2.3.1]. In particular, a similar discussion about the independence of local coordinate charts is contained in [MP25, Proposition 2.3.4].
@@
When restricted to each $U=U_{t}$, the valuation $v_{\ell}$ corresponds to the exceptional divisor $E$ of the weighted blowup of $U_{t}$ along $f(B_{t})\cap U_{t}$ with weights
$(1^{\oplus(d-2)},\ell^{\oplus(n-d+1)})=\left\{\begin{array}[]{ll}(\underbrace{1,\dots,1}_{n-1})&\text{for }\ell=1,\\ (\underbrace{1,\dots,1}_{d-2},\underbrace{2,\cdots,2}_{n-d+1})&\text{for }\ell=2\end{array}\right$.
(See [QR12] for more a general set-up of weighted blowups). In fact, for $\ell=1$ the valuation $v_{1}$ corresponds to the symbolic blowup of $X$ along $Z$.
Fix $\ell\in\{1,2\}$ and set ${\mathcal{I}}_{\lambda}={\mathcal{I}}_{\lambda}(v_{\ell})=\{f\in{\mathcal{O}}_{X};v_{\ell}(f)\geq\lambda\}$. Choose $\lambda=xk$ for $x\in{\mathbb{Q}}_{>0}$. Then similar to the proof of Proposition 3.1, we have the exact sequence:
@@
4.2. Singular minimal rational curves with trivial normal bundles
Our goal in this subsection is to prove the following gap result when there is an immersed rational curve with a trivial bundle and a singular image set.
Proposition 4.2.
Let $X$ be a K-semistable Fano manifold. Assume that $f:{\mathbb{P}}^{1}\rightarrow X$ is an immersed rational curve with a non-empty subset of singular points such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}^{n-1}$. Then we have an estimate of the volume $(-K_{X})^{n}<2n^{n}$. As a consequence, if there exists a minimal rational curve with a trivial normal bundle on $X$ and $(-K_{X})^{n}=2n^{n}$, then the minimal rational curve has no singular points and $X$ must be isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
The rest of this subsection is devoted to proving the above proposition. The last statement follows from the strict volume inequality and Theorem 1.4. So we just need to prove the inequality $(-K_{X})^{n}<2n^{n}$ when $f$ is not an embedding. We use the same notation as the previous subsection but now specialize to the case when $\ell=1$ and $d=2$.
@@ 4.2. Singular minimal rational curves with trivial normal bundles
For any $s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$, choose a local generator $\sigma_{t}$ of $L^{k}$ over $U_{t}$, the section $[s]$ is represented uniquely over $U_{t}$ by a holomorphic function with the following expression (note that, since $d=2$, we get d2=0d-2=0 so that $(z_{1},\dots,z_{n-1},z_{n})=({\mathbf{z}}^{\prime\prime},z_{n})$):
$\frac{s_{t}}{\sigma_{t}}$ $=$ $\sum_{|J|=y}b^{t}_{J}(z_{n}){\mathbf{z}}^{\prime\prime J}$
which we call the initial term representation of $s$ over $U_{t}$. Set
$\gamma_{t}([s])=\sigma_{t}\cdot\sum_{|J|=y}b^{t}_{J}(z_{n})({\rm d}{\mathbf{z}}^{\prime\prime})^{\otimes J}$
which defines a holomorphic section of $L^{k}\otimes\mathrm{Sym}^{y}({\mathcal{O}}^{\oplus(n-1)})$ over $B_{t}$. It is easy to check that $\gamma_{t}([s])$ patches together to define a global section
$\gamma([s])\in H^{0}({\mathbb{P}}^{1},L^{k}\otimes\mathrm{Sym}^{y}({\mathcal{O}}^{\oplus n-1}))=H^{0}({\mathbb{P}}^{1},L^{k})\otimes{\mathbb{C}}^{\binom{n+y-2}{y}}$.
In other words, we know that for each $J=(j_{1},\dots,j_{n-1})$ with $|J|=j_{1}+\cdots+j_{n-1}=y$, $b_{J}^{t}(z_{n})({\rm d}{\mathbf{z}}^{\prime\prime})^{\otimes J}$ patches together to become a section $s_{J}$ in $H^{0}({\mathbb{P}}^{1},L^{k})$.
Now assume that $t\neq t^{\prime}\in{\mathbb{P}}^{1}$ satisfy $f(t)=f(t^{\prime})=x\in f({\mathbb{P}}^{1})=Z\subset X$. Let $C=f(B_{t})$ and $C^{\prime}=f(B_{t^{\prime}})$ be the two distinct locally smooth branches of $Z$ that pass through $x$. Let $(U=U_{t},{\mathbf{z}}={\mathbf{z}}_{t})$ and $(\hat{U}=U_{t^{\prime}},\hat{{\mathbf{z}}}=z_{t^{\prime}})$ be coordinate charts centered at $t$ and $t^{\prime}$ respectively in Subsection 4.1. For each $1\leq k\leq n$, one can write:
$z_{k}$ $=$ $g_{k1}(\hat{z}_{n})\hat{z}_{1}+\cdots+g_{k,{n-1}}(\hat{z}_{n})\hat{z}_{n-1}+\hat{z}_{n}^{m_{k}}h_{k}(\hat{z}_{n})+R_{2},$
where $g_{kj}(\hat{z}_{n})$ and $h_{k}(\hat{z}_{n})$ are holomorphic functions of $\hat{z}_{n}$ with $h_{k}(0)\neq 0$ if $m_{k}>0$, and $R_{2}$ consists of terms of degree at least $2$ in $\hat{z}_{1},\dots,\hat{z}_{n-1}$. Without the loss of generality, assume that $m_{1}\geq m_{2}\geq\cdots\geq m_{n-1}\geq 0$. Then we must have $m_{1}>0$. Otherwise, we have $C^{\prime}=\{\hat{z}_{1}=\cdots=\hat{z}_{n-1}=0\}\subseteq C=\{z_{1}=\cdots=z_{n-1}=0\}$ which is not possible. Assume that $1\leq q\leq n-1$ satisfies $m_{q}>0$ and $m_{q+1}=0$ if q<n1q<n-1. For any $q+1\leq k\leq n-1$, we have ${\rm d}z_{k}|_{T_{x}X}=\sum_{j=1}^{n-1}g_{kj}(0)d\hat{z}_{j}|_{T_{x}X}$ which implies that the matrix $(g_{kj}(0))^{q+1\leq k\leq n-1}_{1\leq j\leq n-1}$ has rank equal to n1qn-1-q. For the section $s_{J}$ obtained from above, we get:
$b^{t}_{J}(z_{n})z_{1}^{j_{1}}z_{2}^{j_{2}}\cdots z_{n-1}^{j_{n-1}}$ $=$ $b^{t}_{J}(z_{n})(\prod_{k=1}^{q}\hat{z}_{k}^{m_{k}}h_{k}(\hat{z}_{n})^{j_{k}}+R_{1})\cdot\prod_{k=q+1}^{n-1}\left(\sum_{j=1}^{n-1}g_{kj}(\hat{z}_{n})\hat{z}_{j}+R_{2}\right)^{j_{k}}$
which should be of the form $b_{J}^{t^{\prime}}(\hat{z}_{n})\hat{z}_{1}^{j_{1}}\hat{z}_{2}^{j_{2}}\cdots\hat{z}_{n-1}^{j_{n-1}}+R_{|J|+1}$, since by construction $f^{*}{\rm d}z_{j}=f^{*}d\hat{z}_{j}$ on $U_{t}\cap U_{t^{\prime}}$ which is the restrction to $U_{t}\cap U_{t^{\prime}}$ of a global non-zero section of the $j$-th factor of ${\mathcal{O}}^{\oplus(n-1)}=N_{f/X}^{\vee}$. This forces $b^{t}_{J}(z_{n})$ to vanish at $z_{n}=0$ up to order at least $j_{1}+\cdots+j_{q}\geq j_{1}$, and moreover $\hat{z}_{n}^{m_{n}}h_{n}(\hat{z}_{n})\equiv 0$. So we see that $s_{J}\in H^{0}({\mathbb{P}}^{1},{\mathcal{O}}(-1)^{j_{1}}\otimes{\mathcal{O}}^{j_{2}}\otimes\cdots\otimes{\mathcal{O}}^{j_{n-1}})$. So we get the estimate:
(17) $h^{0}(C,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})\leq h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes\mathrm{Sym}^{y}({\mathcal{O}}(-1)\oplus{\mathcal{O}}^{(n-2)})).$
By (16) and (17), we get:
(18) $\sum_{y=0}^{xk-1}h^{0}(C,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$ $\leq$ $\sum_{y=0}^{xk-1}h^{0}({\mathbb{P}}^{1},{\mathcal{O}}(2k)\otimes{\rm Sym}^{y}({\mathcal{O}}(-1)\oplus{\mathcal{O}}^{\oplus(n-2)}))$
$\leq$ $\sum_{y=0}^{xk-1}\sum_{i=0}^{y}(2k-i+1)a_{y,i}$
where $a_{y,i}=\binom{n-3+y-i}{n-3}$.
If $x\leqslant 2$, then by using the explicit expression of binomial coefficient, the right-hand-side of (18) is equal to $b_{1}k^{n}/n!+O(k^{n-1})$ where $b_{1}$ is given by the following integral (see (22) for a derivation of a more complicated formula):
$b_{1}$ $=$ $n!\int_{0}^{x}{\rm d}t\int_{0}^{t}(2-z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z=\frac{n!}{(n-2)!}\int_{0}^{x}(2-z)(x-z)^{n-2}{\rm d}z$
$=$ $2nx^{n-1}-x^{n}<2nx^{n-1}.$
When $x>2$, the right-hand-side of (18) splits into the sum: $\sum_{y=0}^{2k}\sum_{i=0}^{y}+\sum_{y=2k+1}^{xk-1}\sum_{i=0}^{2k}$ and the right-hand-side of (18) is equal to $b^{\prime}_{1}k^{n}/n!+O(k^{n-1})$ with
$b^{\prime}_{1}$ $=$ $n!\int_{0}^{2}{\rm d}t\int_{0}^{t}(2-z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z+n!\int_{2}^{x}{\rm d}t\int_{0}^{2}(2-z)\frac{(t-z)^{n-3}}{(n-3)!}{\rm d}z$
$=$ $\frac{n!}{(n-2)!}\int_{0}^{2}(2-z)(x-z)^{n-2}{\rm d}z$
$=$ $2nx^{n-1}-x^{n}+(x-2)^{n}<(2n-1)x^{n-1}+2^{n-1}.$
Set $\psi(x)=2nx^{n-1}$ and
(19) ${\phi}(x)=\left\{\begin{array}[]{ll}2nx^{n-1}&\text{if }\;0\leq x\leq 2\\ (2n-1)x^{n-1}+2^{n-1}&\text{if }\;x\geq 2,\end{array}\right$.
so that $\phi(x)<\psi(x)$ with strict inequality when $x>2$. Moreover it is easy to check that $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ and the strict inequality holds when $V>2n$. So we can apply Lemma 2.4 to conclude that $F^{\phi}(V)\geq F^{\psi}(V)$ and the strict inequality holds for $V>2n$. From Lemma 2.4 and Corollary 2.5, we get $(-K_{X})^{n}\leq(F^{\phi})^{-1}(n-1)<(F^{\psi})^{-1}(n-1)=2n^{n}$ (see Example 2.7). This completes the proof of Proposition 4.2.
4.3. Minimal rational curves with degree $d\geq 3$
In this section, we are going to prove the following result which resolves a large part of Theorem 1.1. Recall the invariant $l_{X}$ from (7), whose definition is the smallest degree of minimal rational curves on $X$.
Proposition 4.3.
Let $X$ be a K-semistable Fano manifold with $3\leq l_{X}\leq n-1$. Then ${\rm vol}(X)<2n^{n}$.
We use the notation from subsection 4.1 and specialize to the case when $\ell=2$ and $3\leq d\leq n+1$.
For any $s]\in H^{0}(X,L^{k}\otimes{\mathcal{I}}_{y}/{\mathcal{I}}_{y+1})$, choose a local generator $\sigma_{t}$ of $L^{k}$ over $U_{t}$, the section $[s]$ is represented uniquely over $U_{t}$ by a holomorphic function with the following expression: